The length of a rectangle is three times its width. If the perimeter of the rectangle is 56 feet, find its area.





Answer by checkley77(12844)

L=3W
P=2L+2W
56=2*3W+2W
56=6W+2W
8W=56
W=56/8
W=7 ANS. FOR THE WIDTH.
L=3*7=21 ANS. FOR THE LENGTH.
PROOF:
56=2*21+2*7
56=42+14
56=56